# Syntactic Language Change in English and German: Metrics, Parsers, and Convergences

### v1

@@ 5.1 Sensitivity of metrics to data noise
While our results in §4 indicate that the parsers perform decently on the target corpora, it remains uncertain whether the metrics derived from their parsing results are affected by data noise and to what extent. Since we will inspect the trends of the metrics over time, it is important to ensure that the data noise does not disrupt the rankings of those metrics. Hence, we compute Spearman’s correlations between our 15 metrics on the original and human corrected sentences (without spelling and OCR errors, etc.; collected in Section 3); higher correlations indicate the rankings are less disrupted and vice versa.
| Metrics | Stanza | CoreNLP | StackPointer | Biaffine | TowerParse | AVG |
| mDD | 0.976 | 0.905 | 0.823 | 0.966 | 0.888 | 0.912 |
| nDD | 0.995 | 0.892 | 0.862 | 0.985 | 0.960 | 0.939 |
| 0pt | 0.983 | 0.915 | 0.862 | 0.986 | 0.820 | 0.913 |
| 0.975 | 0.954 | 0.959 | 0.972 | 0.909 | 0.954 | |
| treeDegree | 0.968 | 0.916 | 0.870 | 0.973 | 0.815 | 0.908 |
| #Leaves | 0.998 | 0.999 | 0.908 | 0.997 | 0.832 | 0.947 |
| degreeVar | 0.966 | 0.935 | 0.923 | 0.976 | 0.834 | 0.927 |
| degreeMean | 0.998 | 0.988 | 0.902 | 0.998 | 0.832 | 0.944 |
| depthVar | 0.982 | 0.910 | 0.861 | 0.976 | 0.827 | 0.911 |
| depthMean | 0.987 | 0.918 | 0.874 | 0.980 | 0.818 | 0.915 |
| 0.989 | 0.967 | 0.891 | 0.990 | 0.841 | 0.936 | |
| #Crossing | 0.942 | - | 0.815 | 0.938 | 0.782 | 0.869 |
| 0.995 | 0.935 | 0.874 | 0.994 | 0.884 | 0.936 | |
| 0.999 | 0.870 | 0.873 | 0.965 | 0.976 | 0.937 | |
| 0.991 | 0.991 | 0.894 | 0.986 | 0.830 | 0.938 | |
| AVG | 0.983 | 0.935 | 0.879 | 0.979 | 0.857 | 0.926 |
Overall, as Figure 5 illustrates, the metrics on the original and corrected sentences strongly correlate with each other, shown with an overall correlation of 0.926. This overall correlation is computed by first averaging over parsers for each metric and then averaging over all metrics. The lowest average correlation per metric is observed with #crossing, 0.869, which, however, still implies a strong correlation. The least sensitive metric is the head-final ratio, shown with an average correlation of 0.954 over parsers. Among the parsers, Stanza is least sensitive (0.983), followed by Biaffine (0.979), whereas TowerParse and StackPointer are the most sensitive ones (0.857 and 0.879 respectively). For this reason, we conclude that our metrics are not much sensitive to data noise in the target corpora across different parsers.
@@ 6.1 What is the dependence of language change trends on parsers?
The outputs of MK can be seen as the labels of a three-way classification task. Therefore, we calculate Cohen’s Kappa to indicate the agreement between each parser pair. The values in Figure 12 are derived by calculating Cohen’s Kappa across the 135 MK outputs based on different parsers for each language. We then average the agreements over all parser pairs containing a certain parser to achieve the average agreement for each parser, shown in the last rows in Figure 12.
@@
Next, we examine the parsers’ agreement on a trend for each metric. We regard the MK result based on each parser as an ‘annotation’ for the corresponding trend; thus, for each trend, we have 5 different ‘annotations’, allowing for calculating the Fleiss’ Kappa (Fleiss, 1971) among them. For each metric, we have 9 ‘annotation’ instances, resulting from the 9 sentence lengths. As Figure 13 shows, most values range from 0.2 to 0.5, implying a consistently low-level of agreement among the parsers across the metrics. treeDegree is the only metric where the parsers obtain a high agreement across both languages (0.8 Fleiss’ Kappa). Among the metrics, parsers least agree on the trends in mDD for English (0.1 Fleiss’ Kappa). This questions the previous works Lei and Wen (2020); Juzek, Krielke, and Teich (2020); Zhang and Zhou (2023) in reporting diachronic trend results based on a single parser, as results may change when using a different parser.
@@ 6.2 Are syntactic changes more alike or distinct in English and German?
We compare the 135 trends between English and German for the 15 metrics and 9 sentence lengths. As shown in Figure 14 (left), among all the trends, 99 cases (73.3%) are incomparable since there are no significant trends for German or English or both; 32 cases (23.7%) have the same trends, while only 4 cases (3%) show different trends, which may suggest a more uniform, rather than diverse, syntactic change between English and German. Among the incomparable trends, as shown in Figure 14 (right), 44 cases (44.4%) are the result of no significant trends for both English and German, 11 cases (11.1%) lack significant trends for German, and 44 cases (44.4%) have no significant trends for English. From this perspective, the German language has undergone more significant syntactic changes compared to English.
@@ 6.6 What are the similarities in syntactic changes between English and German?
For the metrics assessing dependency trees vertically, namely 0pt (Figure 17(a)), depthVar (Figure 17(b)) and depthMean (Figure 17(c)), the trends are similar: they decrease over time for short sentences having 5-7 and/or 10-12 words but increase for long sentences like those of length 60-62 and 70-72. In contrast, the degree-relevant metrics, which assess trees horizontally, increase for short sentences but decrease for long sentences: in Figure 17(e), 17(d) and 17(f), we observe an increase in degreeVar, treeDegree, and degreeMean for sentences having 5-7 and/or 10-12 words, as well as a decrease in both treeDegree and degreeVar for longer sentences of lengths from 40-42 to 60-62 words. All of the above suggests that generally, for both English and German, the dependency trees of shorter sentences with 10-15 words become shorter and wider over time, while those of longer sentences having 30 words become taller and narrower over time. Short and wider tree structures have a higher chance that a word governs multiple dependents compared to the taller and narrower ones and vice versa. An example for comparing between the shorter and wider vs. the taller and narrower dependency trees is given in Figure 19, where both sentences have 7 tokens.
The trends of and are illustrated in Figure 17(g) and 17(i) respectively. A larger of a sentence indicates greater distance from the corresponding fully head-final sentence (where all dependency pairs of that sentence are head-final); thus, it suggests that the sentence is more head-initial. is expected to rise with declined to reflect language becoming more head-initial and vice versa. We see that: (1) decreases for sentences with 5-7 and 15-17 words but increases for those having 70-72 words, implying that short sentences become more head-final, while long sentences become more head-initial; (2) decreases over time for sentences having 30 words, also suggesting a tendency for longer sentences being more head-initial over time.
@@ 7.0.1 Effect of sentence length
As mentioned in §3.2, it is important to control the sentence length for diachronic observations as it may be a confounding variable. Thus, we compute the diachronic trend of the average sentence length per decade group for each length group in our data, visualized in Figure 20. Even if the length difference within one group is up to 3 tokens in our data, we still often see a decrease in average sentence length over time, especially for the longer sentences with more than 50 tokens (last rows in Figure 20). However, the difference in the average sentence length within one length group is only up to 0.1 token, which we deem as an acceptable level. The reasons for this are the following: (1) Intuitively, such a small difference in sentence length is unlikely to influence our metrics substantially. For instance, it takes at least one additional token to increase the maximal possible 0pt and treeDegree. (2) We observe trends of metrics that counter the development of sentence lengths, e.g., that longer sentences have a higher chance to produce taller dependency trees compared to shorter sentences; however, we see uptrends in 0pt in Figure 17(a) for sentences with 60-70 words, whose lengths are overall decreasing over time for both languages. This suggests that such a minor decrease in sentence length does not substantially impact the observations on syntactic changes. (3) The subtrends in metrics do not often follow the subtrends in sentence length. For instance, in 6 out of 8 metrics in Figure 18 for sentences of length 5-7 (0pt, depthVar, depthMean, treeDegree, degreeVar, ), the corresponding trends behave consistently across both languages from the time window 1860-1870 to 1880-1890, even when German sentences become longer while English sentences become shorter during that period.
@@ 7.0.2 Scale of syntactic changes
We note that many of our syntactic trends, even if they are statistically significant, are small in scale. For example, degreeMean in Figure 16(b) changes from slightly below 0.984 to slightly above 0.984 in English over the past 160 years (not all trends are on such low scales, however: decreases from slightly below 14 to slightly above 13 in the same time period in English (Figure 16(c)), for example). On the one hand, this makes sense as syntax is expected to change much slower, e.g., compared to lexical or semantic changes. On the other hand, this may also mean that some of our identified changes may hardly be noticeable (by linguist experts) in actual language or will more pronouncedly manifest themselves only in the future.
@@ 8 Conclusion
In conclusion, this study delved into diachronic language changes in English and German. We comprehensively examined 15 metrics, including the mean dependency distance, across extensive corpora containing political debates. By analyzing the agreement among five dependency parsers, including the commonly used Stanford CoreNLP, we provided novel insights into the importance of used parsers in such language change research. Notably, we discovered a distinctive uptrend in the mean dependency distance in German, which is consistent across various parsers and sentence lengths. We also found that German and English seemed undergo similar syntactic changes in the two political corpora examined, with only few instances showing opposing trends. We also observed more syntactic diachronic variability in sentences of extreme lengths, e.g., below or above mean sentence lengths.
Possible limitations of this work are the following: (1) The text genre explored was restricted to political debates, in order to make the subcorpora comparable, but this may be an important factor affecting the measures. While there are studies showing a trend towards colloquialization in the parliamentary debates of some languages Hiltunen, Räikkönen, and Tyrkkö (2020); Hou and Smith (2021), this has in fact not yet been studied for German. It is possible that one factor in the increase of dependency distance in the present study was in fact one such colloquialization trend, namely that an earlier more nominal style (typical of German written registers, especially legalese, replacing finite clauses with nominal constructions) gave way to a more natural/colloquial ‘verbal’ style with more finite clauses. For instance, a construction like (8) might have been replaced by a construction like (8).
[…]
@@
We will look at this possibility more closely in a future study. (2) The time depth of the used diachronic corpora only covers the last 200 years—which may be too short to draw conclusions about syntactic change, which tends to take several centuries to unfold (e.g. Kroch 2001). (3) We did not thoroughly investigate the origins of the changes in measures. For instance, Juzek, Krielke, and Teich (2020) indicate that the downtrend in dependency distance over time is because dependency relations with shorter distances appear more frequently in the later periods and vice versa, while Liu, Zhu, and Lei (2022) show that longer sentences have more dependency relations with decreasing dependency distances compared to the short ones, which results in the discrepancy in the trends for short and long sentences.

### v2

@@ 5.1 Sensitivity of metrics to data noise
While our results in §4 indicate that the parsers perform decently on the target corpora, it remains uncertain whether the metrics derived from their parsing results are affected by data noise and to what extent. Since we will inspect the trends of the metrics over time, it is important to ensure that the data noise does not disrupt the rankings of those metrics. In §9.3 in the appendix, we demonstrate the insensitivity of our metrics to data noise by showing that the metrics for the clean data are highly correlated to those for the noisy data (overall 0.926 Spearman).
@@ 6.1 What is the dependence of language change trends on parsers?
The outputs of MK can be seen as the labels of a three-way classification task. Therefore, we calculate Cohen’s Kappa (Cohen, 1960) to indicate the agreement between each parser pair. The values in Figure 9 are derived by calculating Cohen’s Kappa across the 135 MK outputs based on different parsers for each language. We then average the agreements over all parser pairs containing a certain parser to achieve the average agreement for each parser, shown in the last rows in Figure 9.
@@
Next, we examine the parsers’ agreement on a trend for each metric. We regard the MK result based on each parser as an ‘annotation’ for the corresponding trend; thus, for each trend, we have 5 different ‘annotations’, allowing for calculating the Fleiss’ Kappa (Fleiss, 1971) among them. For each metric, we have 9 ‘annotation’ instances, resulting from the 9 sentence lengths. As Figure 10 shows, most values range from 0.2 to 0.5, implying a consistently low level of agreement among the parsers across the metrics. treeDegree is the only metric where the parsers obtain a high agreement across both languages (0.8 Fleiss’ Kappa). Among the metrics, parsers show the second least agreement on trends in mDD for English (0.1 Fleiss’ Kappa). This further questions the previous works in reporting diachronic trends of dependency distance based on a single parser, as results may change when using a different parser.
@@ 6.2 Are syntactic changes more alike or distinct in English and German?
We compare the 135 trends between English and German for the 15 metrics and 9 sentence lengths. As shown in Figure 11 (left), among all the trends, 99 cases (73.3%) are incomparable since there are no significant trends for German or English or both; 32 cases (23.7%) have the same trends, while only 4 cases (3%) show different trends, which may suggest a more uniform (“convergence”), rather than diverse, syntactic change between English and German. Among the incomparable trends, as shown in Figure 11 (right), 44 cases (44.4%) are the result of no significant trends for both English and German, 11 cases (11.1%) lack significant trends for German, and 44 cases (44.4%) have no significant trends for English. From this perspective, the German language has undergone more significant syntactic changes compared to English.
@@ 6.6 What are the similarities in syntactic changes between English and German?
For the metrics assessing dependency trees vertically, namely 0pt (Figure 15(a)), depthVar (Figure 15(b)) and depthMean (Figure 15(c)), the trends are similar: they decrease over time for short sentences having 5-7 and/or 10-12 words but increase for long sentences like those of length 60-62 and 70-72. In contrast, the degree-relevant metrics, which assess trees horizontally, increase for short sentences but decrease for long sentences: in Figure 15(e), 15(d) and 15(f), we observe an increase in degreeVar, treeDegree, and degreeMean for sentences having 5-7 and/or 10-12 words, as well as a decrease in both treeDegree and degreeVar for longer sentences of lengths from 40-42 to 60-62 words. All of the above suggests that generally, for both English and German, the dependency trees of shorter sentences with 10-15 words become shorter and wider over time, while those of longer sentences having 30 words become taller and narrower over time. Short and wider tree structures have a higher chance that a word governs multiple dependents compared to the taller and narrower ones and vice versa. An example for comparing between the shorter and wider vs. the taller and narrower dependency trees is given in Figure 16, where both sentences have 7 tokens.
The trends of and are illustrated in Figure 15(g) and 15(i) respectively. A larger of a sentence indicates greater distance from the corresponding fully head-final sentence (where all dependency pairs of that sentence are head-final); thus, it suggests that the sentence is more head-initial. is expected to rise with declined to reflect language becoming more head-initial and vice versa. We see that: (1) decreases for sentences with 5-7 and 15-17 words but increases for those having 70-72 words, implying that short sentences become more head-final, while long sentences become more head-initial; (2) decreases over time for sentences having 30 words, also suggesting a tendency for longer sentences being more head-initial over time.
@@ 7.0.1 Effect of sentence length
As mentioned in §3.2, it is important to control the sentence length for diachronic syntactic observations as it may be a confounding variable.161616On the other hand, reducing sentence length over time may be one core aspect of syntactic simplification, e.g., splitting one long sentence into two shorter sentences. Thus, we compute the diachronic trend of the average sentence length per decade group for each length group in our data, visualized in Figure 17. Even if the length difference within one group is up to 3 tokens in our data, we still often see a decrease in average sentence length over time, especially for the longer sentences with more than 50 tokens (last rows in Figure 17). However, the difference in the average sentence length within one length group is only up to 0.1 token, which we deem as an acceptable level. The reasons for this are the following: (1) Intuitively, such a small difference in sentence length is unlikely to influence our metrics substantially. For instance, it takes at least one additional token to increase the maximal possible 0pt and treeDegree. (2) We observe trends of metrics that counter the development of sentence lengths, e.g., that longer sentences have a higher chance to produce taller dependency trees compared to shorter sentences; however, we see uptrends in 0pt in Figure 15(a) for sentences with 60-70 words, whose lengths are overall decreasing over time for both languages. This suggests that such a minor decrease in sentence length does not substantially impact the observations on syntactic changes. (3) The subtrends in metrics do not often follow the subtrends in sentence length. For instance, in 6 out of 8 metrics in Figure 15 for sentences of length 5-7 (0pt, depthVar, depthMean, treeDegree, degreeVar, ), the corresponding trends behave consistently across both languages from the time window 1860-1870 to 1880-1890, even when German sentences become longer while English sentences become shorter during that period.
@@ 7.0.2 Scale of syntactic changes
We note that several of our syntactic trends, even if they are statistically significant, are small in scale. For example, degreeMean in Figure 14(b) changes from slightly below 0.984 to slightly above 0.984 in English over the past 160 years (not all trends are on such low scales, however: decreases from slightly below 14 to slightly above 13 in the same time period in English (Figure 14(c)), for example). On the one hand, this makes sense as syntax is expected to change much slower, e.g., compared to lexical or semantic changes. On the other hand, this may also mean that some of our identified changes may hardly be noticeable (by linguist experts) in actual language or will more pronouncedly manifest themselves only in the future.
@@ 8 Conclusion
This study delved into diachronic language changes in English and German. We comprehensively examined 15 metrics, including the mean dependency distance, across extensive corpora containing political debates. By analyzing the agreement among five dependency parsers, including the commonly used Stanford CoreNLP, we provided novel insights into the importance of used parsers in such language change research. Notably, we discovered a distinctive uptrend in the mean dependency distance in German, which is consistent across various parsers and sentence lengths. We also found that German and English seemed undergo similar syntactic changes in the two political corpora examined, with only few instances showing opposing trends. We also observed more syntactic diachronic variability in sentences of extreme lengths, e.g., below or above mean sentence lengths. Our results may be relevant in downstream tasks, e.g., for text generation systems that map across historic time frames, e.g., diachronic summarizers or MT systems (Zhang, Ouni, and Eger, 2023; Thai et al., 2022).
Possible limitations of this work are the following: (1) The text genre explored was restricted to political debates, in order to make the subcorpora comparable, but this may be an important factor affecting the measures. While there are studies showing a trend towards colloquialization in the parliamentary debates of some languages Hiltunen, Räikkönen, and Tyrkkö (2020); Hou and Smith (2021), this has in fact not yet been studied for German. It is possible that one factor in the increase of dependency distance in the present study was in fact one such colloquialization trend, namely that an earlier more nominal style (typical of German written registers, especially legalese, replacing finite clauses with nominal constructions) gave way to a more natural/colloquial ‘verbal’ style with more finite clauses. This can be illustrated with the example in (8). (8) is an attested example from the corpus. Such complex nominal structures might over time have been replaced by finite clauses as in the constructed example in (8).
die Beschlüsse, welche das Haus bei §55 durch die Annahme der Kommissionsvorschläge gefaßt hat. (1876-11-21)
@@
(2) The time depth of the used diachronic corpora only covers the last 200 years—which may be too short to draw conclusions about syntactic change, which tends to take several centuries to unfold (e.g. Kroch 2001).
@@
(3) We did not thoroughly investigate the origins of the changes in measures. For instance, Juzek, Krielke, and Teich (2020) indicate that the downtrend in dependency distance over time is because dependency relations with shorter distances appear more frequently in the later periods and vice versa,
while Liu, Zhu, and Lei (2022) show that longer sentences have more dependency relations with decreasing dependency distances compared to the short ones, which results in the discrepancy in the trends for short and long sentences.